OpenStudy (noahbred):

Really need help Which inequality models this problem? The length of a rectangle is five times its width. If the perimeter is at most 96 centimeters, what is the greatest possible value for the width? A. 2w + 2 • (5w) ≤ 96 B. 2w + 2 • (5w) ≥ 96 C. 2w + 2 • (5w) < 96 D. 2w + 2 • (5w) > 96

OpenStudy (noahbred):

@SolomonZelman help?~

OpenStudy (anonymous):

what do u think

OpenStudy (noahbred):

i dont really know i just need help please

OpenStudy (wolf1728):

Okay I'd start with L = 5*w perimeter = 2*L + 2*5*w perimeter = 2L + 10w

OpenStudy (wolf1728):

96 = 2*L + 10w

OpenStudy (noahbred):

Ohhhhhhhh so its greater than or equal right?

OpenStudy (wolf1728):

Actually all 4 answers are in terms of the width. My equation 96 = 2*L + 10w is in terms of width AND length

OpenStudy (noahbred):

sorry i am really dumb id get it still

OpenStudy (wolf1728):

Basically all 4 of the answers are stated in terms of width only. My equation 96 = 2*L + 10w I don't know how they got 2*L to become 2*w This is about as far as I can go.

OpenStudy (noahbred):

oh okay thanks :D

OpenStudy (noahbred):

The length of a rectangle is five times its width. If the perimeter is at most 96 centimeters, what is the greatest possible value for the width? A. 40 cm B. 19.2 cm C. 16 cm D. 8 cm

OpenStudy (noahbred):

Help with this?